OpenStudy (anonymous):

Use mathematical induction to prove that the statement is true for every positive integer n. 2 is a factor of n^2 - n + 2

OpenStudy (anonymous):

n^2-2n+n+2 n*(n-2)+1*(n-2) (n-2)*(n+1) n=2,1 so 2 is the factor

OpenStudy (anonymous):

thank you @speedymath

OpenStudy (anonymous):

First check if the statement is true for n = 0. \[\bf (0)^2-(0)+2=2\]So that works. Now assume that the statement is true for when n = k. Now we must prove the case for n = k + 1. \[\bf (k+1)^2-(k+1)+2=k^2+k+2\]Now it should be noted that if k is even, i.e. k = 2n, then the expression is always true. So now we must prove that the statement is true if k was odd, i.e. k = 2n + 1\[\bf (2n+1)^2+(2n+1)+2=4n^2+6n+4=2(2n^2+3n+2)\]Clearly, we can see that when k is odd, the statement still holds as 2 is a factor of the final expression. Hence 2 is a factor of n^2 - n + 2. @ska8terchick_

OpenStudy (anonymous):

wow thank you @genius12 your name do you justice

OpenStudy (anonymous):

@ska8terchick_ fan/medal? lol -.-